Let A and B be two disjoint events such that P(A) = 0.25 and P(B) = 0.45. What is P(A or B)?
Answer #1

P(Aor B)=P(A)+P(B) = 0.25 +0.45 = 0.70 Therefore, the probability of A or B is 0.70.